According to official census figures, 8% of couples living together ate not married. A reacher took a random sample of 400 couples and found that 9.5% of them are not married. Test at the 15% significance level if the current percentage of unmarried couples is different from 8%
1
Expert's answer
2021-08-06T11:38:57-0400

Because this sample meets the "np=400(0.08)=32>10" and "n(1-p)=400(1-0.08)" "=368>10" conditions, we can assume it is approximately normal. 

The following null and alternative hypotheses for the population proportion needs to be tested:

"H_0: p=0.8"

"H_1: p\not=0.8"

This corresponds to a two-tailed test, for which a z-test for one population proportion will be used.

Based on the information provided, the significance level is "\alpha=0.15," and the critical value for a two-tailed test is "z_c=1.4395."

The rejection region for this two-tailed test is "R=\{z:|z|>1.4395\}."  

The z-statistic is computed as follows:


"z=\dfrac{\hat{p}-p_0}{\sqrt{\dfrac{p_0(1-p_0)}{n}}}=\dfrac{0.095-0.08}{\sqrt{\dfrac{0.08(1-0.08)}{400}}}"

"\approx1.1058"

Since it is observed that "|z|=1.1058<1.4395=z_c," it is then concluded that the null hypothesis is not rejected.

Using the P-value approach: The p-value is "p=2P(Z>1.1058)=0.2688," and since "p=0.2688>0.15=\alpha," it is concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population proportion "p" is different than "0.08," at the "\alpha=0.15" significance level.